JEF meeting, February 28, 2025

From GlueXWiki
Jump to: navigation, search

Connection

Meeting time and location:

Agenda

Preparation of parts for ECAL -- Sasha

  • Waiting for a cable to connect the newly installed MPOD with the light source
  • Parts 3-D printed for the reference PMTs (Hakob is looking on this) , Bobby is preparing a mounting plate and box.
  • Communicated with Adam H. regarding the alpha sources (some paper work was needed). Two sources will be delivered to the Hall.

Installation of the FCAL2 detector & integration to GlueX -- Sasha, All

  • Installing new LMS light source [1]
    • Signal waveforms [2]
  • Installing LMS electronics [3]
  • Remaining tasks:
    • Install reference PMTs: Attach the alpha sources and optical fibers to the PMTs
    • Connect the PMTs to the HV supply and the readout crate (cables have already been installed)
    • Test the LMS CSS control

Preparing for the run -- Sasha

  • Stability of the ECAL performance. Testing and addressing issues with the ECAL modules
    • ECAL modules with loose PMT connections [4] [5]
  • Calibration and stabilization of the light and humidity sensors [6]
  • Integrating ECAL into the GlueX trigger [7]

Restoring PS test setup [8]

  • Preparing a prototype with 5x5 ECAL modules (all modules exist) and 5x5 LG modules (can use dividers with the 'conventional' power supply)
    • Check if we can refurbish the beam profiler to place in front of the prototype.

ECAL commissioning -- Sasha, All

Calibrations

Minutes

Attendees: Sasha, Vlad, Zisis, Mac, Bill, Igal, Hakob, Liping, Xiaolin, Jane, Olga, S.T.

  • Waiting for cable to connect MPOD to new LMS light source
  • Parts for reference PMT assembly 3D-printed
    • Still need enclosure for radioactive source stack
    • Working on paperwork for radioactive source
    • Reference PMT needs to be incorporated into CSS
  • New light source for LMS with 25 LEDs much brighter
    • Pulse shapes narrower than previous incarnation
    • Used to adjust delays for trigger
    • Some small amplitudes seen with new light source, but these were confirmed to be due to the fibers
  • Checked ECAL stability -- looking for loose connections after maintenance in dark box
    • Another LMS scan found 7 loose connections, now fixed after 1.5 days of effort
  • Some issues with dark room sensors deviating from previous measurements, probably due to power work in dark room.
  • Working on protocol for working in dark room
    • When going in for maintenance, turn off fans as well as HV
  • Working on integration of ECAL into trigger
    • Will be separate component in energy sum
    • New GTP firmware installed
    • Checked VTP firmware
    • New DAQ code to be incorporated into main DAQ system
  • Planning to use PS arm in Hall-D to do shower shape studies and look at ECAL-FCAL transition region
    • Profile for y-position
  • Sasha worked on making Zhikun's scripts for setting new HV settings more modular
  • Need to decide which clusterization code to use for reconstruction before REST production